40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 20 00:59:02 2011, kiwamu, Summary, General, Weekly report 
    Reply  Mon Jun 27 10:04:29 2011, kiwamu, Summary, General, Weekly report 
       Reply  Mon Jul 4 14:27:35 2011, kiwamu, Summary, General, Weekly report 
          Reply  Mon Jul 25 11:02:19 2011, kiwamu, Summary, General, Weekly report 
             Reply  Mon Jul 25 11:04:22 2011, kiwamu, Summary, General, Weekly report 
                Reply  Mon Jul 25 11:06:38 2011, kiwamu, Summary, General, Weekly report 
                   Reply  Mon Aug 1 22:30:59 2011, kiwamu, Summary, General, Weekly report 
Message ID: 5028     Entry time: Mon Jul 25 11:06:38 2011     In reply to: 5027     Reply to this: 5085
Author: kiwamu 
Type: Summary 
Category: General 
Subject: Weekly report 

Summary of the week ending July 24th.  Number of elog entries = 45


- LSC
 * Check of LSC WF switching
  -> some were switching, but the majority were not.

- SUS
 * Ran activateDQ.py for seeting some DQ channels of Oplevs
 * Turned ON all the offset buttons on the OL1, etc.
 * Rebuilt and restarted c1msc, c1sus, c1scx and c1scy as an update.
 * ETMY's shadow sensors look bad. Unknown noise below 3 Hz, which is higher than the usual floor by factor of 10.

- ABSL
 * The frequency lock was down.
 * The laser power into the RFPD had been too big, so it was reduced

- OAF
 * Seismometers were connected to the AA-board on 1X6
 * Most of the channels were acquired to the ADC, but some were not.

- Mode Cleaner
 * Gain of quadrants were checked.
 * Due to the SUS model update, the MC locking trigger hasn't worked correctly. This was fixed by changing ioo.db file

- Misc.
 * Virtual box was installed on Rossa. Altium is now available on Rossa.

ELOG V3.1.3-